# A Comprehensive Survey of Scientific Large Language Models and Their Applications in Scientific Discovery

### v1

@@ Abstract
In many scientific fields, large language models (LLMs) have revolutionized the way with which text and other modalities of data (e.g., molecules and proteins) are dealt, achieving superior performance in various applications and augmenting the scientific discovery process. Nevertheless, previous surveys on scientific LLMs often concentrate on one to two fields or a single modality. In this paper, we aim to provide a more holistic view of the research landscape by unveiling cross-field and cross-modal connections between scientific LLMs regarding their architectures and pre-training techniques. To this end, we comprehensively survey over 250 scientific LLMs, discuss their commonalities and differences, as well as summarize pre-training datasets and evaluation tasks for each field and modality. Moreover, we investigate how LLMs have been deployed to benefit scientific discovery. Resources related to this survey are available at https://github.com/yuzhimanhua/Awesome-Scientific-Language-Models. * Equal contribution
@@ 1 Introduction
The emergence of large language models (LLMs) Zhao et al. (2023c) brings a new paradigm to natural language processing (NLP) by replacing specialized models designed for each task with unified models that are reasonably effective for a wide spectrum of problems. In the scientific domain, such a paradigm not only reshapes people’s strategies to handle tasks related to natural language (e.g., scientific papers, medical records, and climate reports) but also inspires analogous ideas to deal with other types of data (e.g., molecules, proteins, tables, and metadata). In addition to understanding existing scientific data, LLMs have shown their potential to accelerate scientific discovery Wang et al. (2023c); Zhang et al. (2023f); Wang et al. (2024b) through generation, planning, etc.
@@
Given the broad and profound impact of LLMs in various scientific fields across diverse modalities, it becomes necessary to comprehensively review related work in this direction. However, existing scientific LLM surveys typically focus on either one to two fields (e.g., biomedicine Wang et al. (2023a); He et al. (2024); Pei et al. (2024) and chemistry Xia et al. (2023); Zhang et al. (2024c)) or one modality (e.g., text Ho et al. (2024)) only. In fact, if we take a holistic view of the research landscape, we can observe similar and interrelated techniques used to develop LLMs for different fields and modalities.
@@
Figure 1 depicts three major types of scientific LLM pre-training strategies (i.e., Columns 1 to 3), for each of which we give 4 examples (i.e., Types a to d). In Column 1, following BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019), existing studies use masked language modeling (MLM) to pre-train encoder language models. Here, the input can be naturally sequential (e.g., papers in each field, protein/DNA/RNA sequences in the FASTA format Lipman and Pearson (1985)) or artificially linearized (e.g., molecules in the SMILES format Weininger (1988), sequences of venue/author/reference nodes in citation graphs). In Column 2, inspired by GPT Brown et al. (2020) and LLaMA Touvron et al. (2023a), previous studies adopt next token prediction to pre-train (encoder-)decoder language models, some of which further adopt instruction tuning and preference optimization Ouyang et al. (2022). Other than plain text input (e.g., question-answer pairs from knowledge bases or exams), we see more ways to sequentialize complex scientific data, such as flattening table cells and using particle coordinates to describe crystals. Even for images, there are studies in both mathematics Gao et al. (2023) and biomedicine Li et al. (2023a) that exploit a vision encoder to project an image onto several visual tokens and prepend them to text tokens as linearized LLM input. In Column 3, following DPR Karpukhin et al. (2020) and CLIP Radford et al. (2021), two encoders are pre-trained to map relevant data pairs closer in the latent space via contrastive learning. When both modalities are sequential (e.g., text-text or text-protein), the model is built upon two LLM encoders. When we prefer to keep the non-sequential nature of one modality (e.g., molecular graphs Edwards et al. (2021), chest X-rays Zhang et al. (2022), and aerial views Yan et al. (2024)), the corresponding graph or image encoder can be employed. To summarize, a cross-field cross-modal survey will more accurately draw the connections between different scientific LLMs, demonstrate their commonalities, and potentially guide their future designs.
@@
Contributions. In this paper, motivated by the discussions above, we systematically survey over 250 scientific LLMs encompassing various fields (e.g., general science, mathematics, physics, chemistry, materials science, biology, medicine, and geoscience), modalities (e.g., language, graph, vision, table, molecule, protein, genome, and climate time series), and sizes (from $\sim$100M to $\sim$100B parameters). For each field/modality, we investigate commonly adopted pre-training datasets, model architectures, and evaluation tasks of scientific LLMs. Following our motivation, when we talk about model architectures in detail, we link them back to Figure 1 to build cross-field cross-modal connections. Moreover, we provide a structured summary of these scientific LLMs in Table A1-Table A6 (Appendix A). Furthermore, for different fields, we introduce how LLMs have been deployed to benefit science by augmenting different aspects and stages of the scientific discovery process, such as hypothesis generation, theorem proving, experiment design, drug discovery, and weather forecasting.
@@ 2.1 Language
The most commonly used pre-training corpora for scientific LLMs are research papers from bibliographic databases, such as AMiner Tang et al. (2008), Microsoft Academic Graph (MAG) Sinha et al. (2015), and Semantic Scholar Ammar et al. (2018). Some of these sources (e.g., S2ORC Lo et al. (2020)) contain paper full-text information, while the others have titles and abstracts only.
The evolution of scientific LLMs bears similarity to that of general-domain LLMs. Specifically, pioneering models utilize paper text in a self-supervised way during pre-training, aiming to acquire scientific knowledge from large-scale unlabeled corpora. For example, masked language modeling (MLM) is the default pre-training task for scientific LLMs with a BERT backbone (Type 1.a in Figure 1, e.g., SciBERT Beltagy et al. (2019)); next token prediction is widely used for GPT-based scientific LLMs (Type 2.a in Figure 1, e.g., SciGPT Luu et al. (2021)). More recently, inspired by the fact that LLMs can be trained to follow natural language instructions Wei et al. (2022a); Ouyang et al. (2022), researchers have put more effort into tuning LLMs with instructions to solve complex scientific problems (Type 2.a, e.g., Galactica Taylor et al. (2022) and SciGLM Zhang et al. (2024a)). The instruction tuning data are often derived from datasets for downstream tasks, such as exam question answering Welbl et al. (2017), and further filtered/augmented by humans or existing LLMs (e.g., GPT-4 Achiam et al. (2023)).
General scientific LLMs are usually evaluated on common NLP tasks, such as named entity recognition (NER), relation extraction (RE) Luan et al. (2018), question answering (QA) Wang et al. (2023g), and classification Cohan et al. (2019).
@@ 2.2 Language + Graph
Beyond plain text, scientific papers are associated with rich metadata including venues, authors, and references Zhang et al. (2023h). Such metadata connect papers into a graph that complements text signals for characterizing paper semantics. To exploit metadata, some studies (Type 1.b, e.g., OAG-BERT Liu et al. (2022b)) concatenate paper text with venues/authors as input and perform MLM on both text and metadata; others (Type 3.a, e.g., SPECTER Cohan et al. (2020)) take citation links as supervision and train LLMs to encode linked papers closer in the embedding space. Recent approaches further modify the Transformer architecture in LLMs with Adapters Singh et al. (2023), GNN-nested Transformers Jin et al. (2023b), and Mixture-of-Experts Transformers Zhang et al. (2023g) to better capture graph signals.
@@ 2.3 Applications in Scientific Discovery
Performant scientific LLMs can work alongside researchers throughout the entire scientific discovery process. Leaving field-specific applications for later sections, here we underscore LLMs’ general usefulness in brainstorming and evaluation: Lahav et al. (2022) integrate LLMs into a search engine for the discovery of scientific challenges and directions; Wang et al. (2023f) and Baek et al. (2024) leverage LLMs to generate novel scientific ideas grounded in prior literature; Zhang et al. (2023i) rely on LLMs to find expert reviewers for each submission; Liu and Shah (2023), Liang et al. (2023a), and D’Arcy et al. (2024) explore the capacity of GPT-4 to provide useful feedback on research papers to facilitate automatic review generation.
@@ 3.1 Language
The pre-training text corpora for math LLMs can be categorized into two classes: (1) multiple-choice QA, the representative datasets of which include MathQA Amini et al. (2019), Ape210K Zhao et al. (2020), and Math23K Wang et al. (2017); as well as (2) generative QA, the representative datasets of which include GSM8K Cobbe et al. (2021a), MATH Hendrycks et al. (2021), and MetaMathQA Yu et al. (2023b).
Similarly to general science LLMs, the backbone model of pioneering math LLMs is BERT (Type 1.a, e.g., GenBERT Geva et al. (2020) and MathBERT Shen et al. (2021)). These models are mostly trained via MLM, with the only exception being BERT-TD Li et al. (2022c), where a contrastive loss is adopted. For GPT-based math LLMs (Type 2.a, e.g., GSM8K-GPT Cobbe et al. (2021b) and NaturalProver Welleck et al. (2022)), we find a diversity of pre-training tasks: supervised fine-tuning, next token prediction, and instruction tuning. The most recent math LLMs (Type 2.a, e.g., Rho-Math Lin et al. (2024a) and MAmmoTH2 Yue et al. (2024)) are based on LLaMA and are trained to follow natural language instructions. However, when an enormous pre-training dataset (e.g., 55 billion tokens) is available, next token prediction is still favored as the mere pre-training task Azerbayev et al. (2023); Lin et al. (2024a) or the companion task Shao et al. (2024); Ying et al. (2024) to build base models.
QA and math world problems (MWP) have been the most common evaluation tasks. In addition, quantitative reasoning contains more difficult problems, as the model has to provide a complete and self-contained solution without relying on external tools Shao et al. (2024); Lin et al. (2024a). We see a dominance of use from GSM8K and MATH for QA, and from MathQA and Math23K for MWP. For quantitative reasoning, MMLU-STEM Hendrycks et al. (2020) and Big-Bench Hard Suzgun et al. (2023) are the most widely adopted.
@@ 3.2 Language + Vision
The key to incorporating the vision modality to LLMs is to encode the images and obtain linearized visual representations. Specifically, Inter-GPS Lu et al. (2021) (Type 2.d) uses RetinaNet Lin et al. (2017) to transform images into a set of relationships and then applies BART Lewis et al. (2020a) to produce the solution; G-LLaVA Gao et al. (2023) (Type 2.d) encodes visual input via a pre-trained vision Transformer (ViT), concatenates visual embeddings with textual embeddings, and then feeds the concatenation into LLaMA-2 Touvron et al. (2023b). These models are by default pre-trained via sequence-to-sequence tasks, where the problem is the input, and the ground-truth answer with optional rationale is the output. Auxiliary loss such as masked image modeling, image construction, or text-image matching, is optionally added for better visual modeling.
@@ 3.3 Table
The challenge in tables is similar to that in diagrams, namely to obtain linearized table representations. In most cases, tables are squeezed into linear text sequences as part of the context and are prepended with the question text as the model input. As one of the first works in this line of research, TAPAS Herzig et al. (2020) (Type 1.a) adopts the MLM object to predict the masked token in textual and tabular context. The most recent developments Li et al. (2023c); Zhang et al. (2024d) resemble the design of TableLlama Zhang et al. (2023d) (Type 2.b), with LLaMA-2 as the backbone and instruction tuning as the pre-training task.
Table LLMs are validated through table QA, where the model is asked to produce the correct answer given the table structure, data values, and a question text. Most existing studies have been evaluated on the WikiTableQuestions and WikiSQL datasets. TableInstruct Zhang et al. (2023d) is the most recently developed comprehensive benchmark integrating 14 datasets across 11 tasks.
@@ 4 LLMs in Physics (Table A3)
Existing physics LLMs largely focus on astronomy and the “Language” modality. As a derivative of BERT, astroBERT Grezes et al. (2021) (Type 1.a) is further pre-trained using astronomy-related papers via MLM and next sentence prediction. It is evaluated on the NER task. Likewise, AstroLLaMA Nguyen et al. (2023b) (Type 2.a) fine-tunes LLaMA-2 using over 300,000 astronomy abstracts from arXiv. It is evaluated on paper generation and paper recommendation tasks. AstroLLaMA-chat Perkowski et al. (2024) (Type 2.a) is the chat version of AstroLLaMA. It is continually trained on a GPT-4 generated domain-specific dialogue dataset.
@@ 5.1 Language
LLM pre-training corpora in chemistry and materials science typically come from research papers and databases (e.g., Materials Project Jain et al. (2013)). Besides, recent works adopt domain-specific instruction tuning datasets (e.g., Mol-Instructions Fang et al. (2023a) and SMolInstruct Yu et al. (2024)) derived from PubChem Kim et al. (2019), MoleculeNet Wu et al. (2018), etc.
Early studies on chemistry LLMs mostly adopt a moderate-sized encoder-only architecture pre-trained with MLM (Type 1.a, e.g., ChemBERT Guo et al. (2022), MatSciBERT Gupta et al. (2022), and BatteryBERT Huang and Cole (2022)). These models are usually evaluated on downstream tasks including reaction role labeling Guo et al. (2022) and abstract classification Gupta et al. (2022). Recently, researchers have focused more on large-scale decoder-only LLMs trained with next token prediction and instruction tuning (Type 2.a). Examples include ChemDFM Zhao et al. (2024), ChemLLM Zhang et al. (2024b), and LlaSMol Yu et al. (2024). Given the desired generalization capability of such models, they are evaluated on a diverse set of tasks such as name conversion Kim et al. (2019), reaction prediction Jin et al. (2017), retrosynthesis Schneider et al. (2016), text-based molecule design Edwards et al. (2022), and crystal generation Antunes et al. (2023); Flam-Shepherd and Aspuru-Guzik (2023); Gruver et al. (2024).

### v2

@@ Abstract
In many scientific fields, large language models (LLMs) have revolutionized the way text and other modalities of data (e.g., molecules and proteins) are handled, achieving superior performance in various applications and augmenting the scientific discovery process. Nevertheless, previous surveys on scientific LLMs often concentrate on one or two fields or a single modality. In this paper, we aim to provide a more holistic view of the research landscape by unveiling cross-field and cross-modal connections between scientific LLMs regarding their architectures and pre-training techniques. To this end, we comprehensively survey over 250 scientific LLMs, discuss their commonalities and differences, as well as summarize pre-training datasets and evaluation tasks for each field and modality. Moreover, we investigate how LLMs have been deployed to benefit scientific discovery. Resources related to this survey are available at https://github.com/yuzhimanhua/Awesome-Scientific-Language-Models. * Equal contribution
@@ 1 Introduction
The emergence of large language models (LLMs) Zhao et al. (2023c) brings a new paradigm to natural language processing (NLP) by replacing specialized models designed for each task with unified models that are reasonably effective for a wide spectrum of problems. In the scientific domain, such a paradigm not only reshapes people’s strategies to handle tasks related to natural language (e.g., scientific papers, medical records, and climate reports) but also inspires analogous ideas to deal with other types of data (e.g., molecules, proteins, tables, and metadata). In addition to understanding existing scientific data, LLMs have shown their potential to accelerate scientific discovery Wang et al. (2023c); Zhang et al. (2023e); Wang et al. (2024d) through generation, planning, etc.
@@
Given the broad and profound impact of LLMs in various scientific fields across diverse modalities, it becomes necessary to comprehensively review related work in this direction. However, existing scientific LLM surveys typically focus on either one or two fields (e.g., biomedicine Wang et al. (2023a); He et al. (2024); Pei et al. (2024); Zhang et al. (2024d) and chemistry Xia et al. (2023); Pei et al. (2024); Zhang et al. (2024d)) or one modality (e.g., text Ho et al. (2024)) only. In fact, if we take a holistic view of the research landscape, we can observe similar and interrelated techniques used to develop LLMs for different fields and modalities.
@@
Figure 1 depicts three major types of scientific LLM pre-training strategies (i.e., Columns 1 to 3), for each of which we give 4 examples (i.e., Types a to d). In Column 1, following BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019), existing studies use masked language modeling (MLM) to pre-train encoder language models. Here, the input can be naturally sequential (e.g., papers in each field, protein/DNA/RNA sequences in the FASTA format Lipman and Pearson (1985)) or artificially linearized (e.g., molecules in the SMILES format Weininger (1988), sequences of venue/author/reference nodes in citation graphs). In Column 2, inspired by GPT Brown et al. (2020) and LLaMA Touvron et al. (2023a), previous studies adopt next token prediction to pre-train (encoder-)decoder language models, some of which further adopt instruction tuning and preference optimization Ouyang et al. (2022). Other than plain text input (e.g., question-answer pairs from knowledge bases or exams), we see more ways to sequentialize complex scientific data, such as flattening table cells and using particle coordinates to describe crystals. Even for images, there are studies in both mathematics Gao et al. (2023) and biomedicine Li et al. (2023a) that exploit a vision encoder to project an image onto several visual tokens and prepend them to text tokens as linearized LLM input. In Column 3, following DPR Karpukhin et al. (2020) and CLIP Radford et al. (2021), two encoders are pre-trained to map relevant data pairs closer in the latent space via contrastive learning. When both modalities are sequential (e.g., text-text or text-protein), the model is built upon two LLM encoders. When we prefer to keep the non-sequential nature of one modality (e.g., molecular graphs Edwards et al. (2021), chest X-rays Zhang et al. (2022), and aerial views Yan et al. (2024)), the corresponding graph or image encoder can be employed. To summarize, a cross-field, cross-modal survey will more accurately draw the connections between different scientific LLMs, demonstrate their commonalities, and potentially guide their future designs.
@@
Contributions. In this paper, motivated by the discussions above, we systematically survey over 250 scientific LLMs encompassing various fields (e.g., general science, mathematics, physics, chemistry, materials science, biology, medicine, and geoscience), modalities (e.g., language, graph, vision, table, molecule, protein, genome, and climate time series), and sizes (from $\sim$100M to $\sim$100B parameters). For each field/modality, we investigate commonly adopted pre-training datasets, model architectures, and evaluation tasks of scientific LLMs. Following our motivation, when we discuss model architectures in detail, we link them back to Figure 1 to build cross-field cross-modal connections. Moreover, we provide a structured summary of these scientific LLMs in Table A1-Table A6 (Appendix A). Furthermore, for different fields, we introduce how LLMs have been deployed to benefit science by augmenting different aspects and stages of the scientific discovery process, such as hypothesis generation, theorem proving, experiment design, drug discovery, and weather forecasting.
@@ 2.1 Language
The most commonly used pre-training corpora for scientific LLMs are research papers from bibliographic databases, such as AMiner Tang et al. (2008), Microsoft Academic Graph (MAG) Sinha et al. (2015), and Semantic Scholar Ammar et al. (2018). Some of these sources (e.g., S2ORC Lo et al. (2020)) contain full-text information of papers, while the others have titles and abstracts only.
The evolution of scientific LLMs bears similarity to that of general-domain LLMs. Specifically, pioneering models utilize paper text in a self-supervised manner during pre-training, aiming to acquire scientific knowledge from large-scale unlabeled corpora. For example, masked language modeling (MLM) is the default pre-training task for scientific LLMs with a BERT backbone (Type 1.a in Figure 1, e.g., SciBERT Beltagy et al. (2019)); next token prediction is widely used for GPT-based scientific LLMs (Type 2.a in Figure 1, e.g., SciGPT Luu et al. (2021)). More recently, inspired by the fact that LLMs can be trained to follow natural language instructions Wei et al. (2022a); Ouyang et al. (2022), researchers have put more effort into tuning LLMs with instructions to solve complex scientific problems (Type 2.a, e.g., Galactica Taylor et al. (2022) and SciGLM Zhang et al. (2024a)). The instruction tuning data are often derived from datasets for downstream tasks, such as exam question answering Welbl et al. (2017), and further filtered or augmented by humans or existing LLMs (e.g., GPT-4 Achiam et al. (2023)).
General scientific LLMs are usually evaluated on common NLP tasks, such as named entity recognition (NER), relation extraction (RE) Luan et al. (2018), question answering (QA) Wang et al. (2024e), and classification Cohan et al. (2019).
@@ 2.2 Language + Graph
Beyond plain text, scientific papers are associated with rich metadata including venues, authors, and references Zhang et al. (2023g). Such metadata connect papers into a graph that complements text signals for characterizing paper semantics. To exploit metadata, some studies (Type 1.b, e.g., OAG-BERT Liu et al. (2022b)) concatenate paper text with venues/authors as input and perform MLM on both text and metadata; others (Type 3.a, e.g., SPECTER Cohan et al. (2020)) take citation links as supervision and train LLMs to encode linked papers closer in the embedding space. Recent approaches further modify the Transformer architecture in LLMs with Adapters Singh et al. (2023), GNN-nested Transformers Jin et al. (2023b), and Mixture-of-Experts Transformers Zhang et al. (2023f) to better capture graph signals.
@@ 2.3 Applications in Scientific Discovery
Performant scientific LLMs can work alongside researchers throughout the entire scientific discovery process. Leaving field-specific applications for later sections, here we underscore LLMs’ general usefulness in brainstorming and evaluation: Lahav et al. (2022) integrate LLMs into a search engine for the discovery of scientific challenges and directions; Wang et al. (2023e), Yang et al. (2024d), Baek et al. (2024), and Gu and Krenn (2024) leverage LLMs to generate novel scientific ideas, directions, and hypotheses on the basis of prior literature and existing knowledge; Zhang et al. (2023h) rely on LLMs to find expert reviewers for each submission; Liu and Shah (2023), Liang et al. (2024c), and D’Arcy et al. (2024) explore the capacity of GPT-4 to provide useful feedback on research papers to facilitate automatic review generation; Liang et al. (2024b, a) also observe the increasing use of LLMs in writing scientific papers and conference peer reviews.
@@ 3.1 Language
The pre-training text corpora for math LLMs can be categorized into two classes: (1) multiple-choice QA, the representative datasets of which include MathQA Amini et al. (2019), Ape210K Zhao et al. (2020), and Math23K Wang et al. (2017); as well as (2) generative QA, the representative datasets of which include GSM8K Cobbe et al. (2021a), MATH Hendrycks et al. (2021), and MetaMathQA Yu et al. (2024c).
Similarly to general science LLMs, the backbone model of pioneering math LLMs is BERT (Type 1.a, e.g., GenBERT Geva et al. (2020) and MathBERT Shen et al. (2021)). These models are mostly trained via MLM, with the only exception being BERT-TD Li et al. (2022c), where a contrastive loss is adopted. For GPT-based math LLMs (Type 2.a, e.g., GSM8K-GPT Cobbe et al. (2021b) and NaturalProver Welleck et al. (2022)), we find a diversity of pre-training tasks: supervised fine-tuning, next token prediction, and instruction tuning. The most recent math LLMs (Type 2.a, e.g., Rho-Math Lin et al. (2024b) and MAmmoTH2 Yue et al. (2024c)) are based on LLaMA and are trained to follow natural language instructions. However, when an enormous pre-training dataset (e.g., 55 billion tokens) is available, next token prediction is still favored as the mere pre-training task Azerbayev et al. (2024); Lin et al. (2024b) or the companion task Shao et al. (2024); Ying et al. (2024) to build base models.
QA and math world problems (MWP) have been the most common evaluation tasks. In addition, quantitative reasoning contains more difficult problems, as the model has to provide a complete and self-contained solution without relying on external tools Shao et al. (2024); Lin et al. (2024b). We see a dominance of use from GSM8K and MATH for QA, and from MathQA and Math23K for MWP. For quantitative reasoning, MMLU-STEM Hendrycks et al. (2020) and Big-Bench Hard Suzgun et al. (2023) are the most widely adopted.
@@ 3.2 Language + Vision
The key to incorporating the vision modality into LLMs is to encode the images and obtain linearized visual representations. Specifically, Inter-GPS Lu et al. (2021) (Type 2.d) uses RetinaNet Lin et al. (2017) to transform images into a set of relationships and then applies BART Lewis et al. (2020a) to produce the solution; G-LLaVA Gao et al. (2023) (Type 2.d) encodes visual input via a pre-trained vision Transformer (ViT), concatenates visual embeddings with textual embeddings, and then feeds the concatenation into LLaMA-2 Touvron et al. (2023b). These models are by default pre-trained via sequence-to-sequence tasks, where the problem is the input, and the ground-truth answer with optional rationale is the output. Auxiliary loss such as masked image modeling, image construction, or text-image matching, is optionally added for better visual modeling.
@@ 3.3 Table
The challenge in tables is similar to that in diagrams, namely to obtain linearized table representations. In most cases, tables are squeezed into linear text sequences as part of the context and are prepended with the question text as the model input. As one of the first works in this line of research, TAPAS Herzig et al. (2020) (Type 1.a) adopts the MLM objective to predict the masked token in textual and tabular context. The most recent developments Li et al. (2024b); Zhang et al. (2024f) resemble the design of TableLlama Zhang et al. (2024e) (Type 2.b), with LLaMA-2 as the backbone and instruction tuning as the pre-training task.
Table LLMs are validated through table QA, where the model is asked to produce the correct answer given the table structure, data values, and a question text. Most existing studies have been evaluated on the WikiTableQuestions and WikiSQL datasets. TableInstruct Zhang et al. (2024e) is the most recently developed comprehensive benchmark integrating 14 datasets across 11 tasks.
@@ 4 LLMs in Physics (Table A3)
4.1 Language
As a derivative of BERT, astroBERT Grezes et al. (2021) (Type 1.a) is further pre-trained using astronomy-related papers via MLM and next sentence prediction. It is evaluated on the NER task. Likewise, AstroLLaMA Nguyen et al. (2023b) (Type 2.a) fine-tunes LLaMA-2 using over 300,000 astronomy abstracts from arXiv. It is evaluated on paper generation and recommendation tasks. AstroLLaMA-chat Perkowski et al. (2024) (Type 2.a) is the chat version of AstroLLaMA. It is continually trained on a GPT-4 generated domain-specific dialogue dataset. PhysBERT Hellert et al. (2024) (Type 1.a) is the first physics-specific model for sentence embedding trained on a curated corpus of physics literature based on 1.2 million physics papers on arXiv. It is evaluated on physics-tailored tasks, such as information retrieval, classification, and semantic similarity estimation.
4.2 Applications in Scientific Discovery
Transformer-based physics LLMs can potentially assist humans in solving differential equations and designing experiments. For instance, Cai et al. (2024) apply Transformer to predict the integer coefficients in the scattering amplitudes of Planar $\mathcal{N}=4$ Super Yang-Mills theory; RydbergGPT Fitzek et al. (2024) uses Transformer to learn the distribution of qubit measurement outcomes that describe an array of interacting Rydberg atoms; Arlt et al. (2024) present an initial trial that applies a code-generating LLM to synthesize experimental blueprints for a whole class of quantum systems in the form of Python code.
@@ 5.1 Language
LLM pre-training corpora in chemistry and materials science typically come from research papers and databases (e.g., Materials Project Jain et al. (2013)). Besides, recent works adopt domain-specific instruction tuning datasets (e.g., Mol-Instructions Fang et al. (2023) and SMolInstruct Yu et al. (2024a)) derived from PubChem Kim et al. (2019), MoleculeNet Wu et al. (2018), etc.
Early studies on chemistry LLMs mostly adopt a moderate-sized encoder-only architecture pre-trained with MLM (Type 1.a, e.g., ChemBERT Guo et al. (2022), MatSciBERT Gupta et al. (2022), and BatteryBERT Huang and Cole (2022)). These models are usually evaluated on downstream tasks including reaction role labeling Guo et al. (2022) and abstract classification Gupta et al. (2022). Recently, researchers have focused more on large-scale decoder-only LLMs trained with next token prediction and instruction tuning (Type 2.a). Examples include ChemDFM Zhao et al. (2024), ChemLLM Zhang et al. (2024b), and LlaSMol Yu et al. (2024a). Given the desired generalization capability of such models, they are evaluated on a diverse set of tasks such as name conversion Kim et al. (2019), reaction prediction Jin et al. (2017), retrosynthesis Schneider et al. (2016), text-based molecule design Edwards et al. (2022), and crystal generation Antunes et al. (2023); Flam-Shepherd and Aspuru-Guzik (2023); Gruver et al. (2024).